# An Empirical Study of Testing Practices in Open Source AI Agent Frameworks and Agentic Applications

### v2

@@ 5.3.3 Closed Card Sorting for Components
With the SUTs identified in the previous step (Section 5.3.2), we performed a closed card-sorting to uncover the higher-level architectural components of agent frameworks and applications being tested. We considered the components described in Section 3.3 as the source of components and grounded each SUT to the relevant component through the closed card sorting process.
@@
6.1 Motivation
As established in the Introduction, the prevailing benchmark-driven evaluation culture often fails to capture the practical reliability of agentic systems, overlooking critical failure modes like prompt decay (Bhargava et al., 2024) and tool-use errors that are already documented in the field (Ma et al., 2024; Kokane et al., 2025). While software testing offers a path to bridge this reliability gap, the inherent non-determinism and non-reproducibility of foundation models (FMs) fundamentally challenge the application of traditional, deterministic testing techniques (Hassan et al., 2024).
This conflict creates a crucial knowledge vacuum: with no established best practices for this new paradigm, it remains unknown how practitioners are adapting their testing strategies to cope with these challenges. Therefore, to bridge this gap, we conduct the first large-scale empirical investigation to systematically identify and catalog the testing patterns that have emerged in practice. We specifically analyze how developers structure their tests to manage variability (the Arrange phase of a test) and what strategies they employ to verify unpredictable yet potentially correct outputs (the Assert phase), as more than 75% of tests follow the AAA pattern in recent times (Wei et al., 2025) in traditional software engineering. The goal is to produce a foundational catalog of these real-world strategies, providing the first empirical baseline on how the open-source community is negotiating the unique quality assurance hurdles of the agentic paradigm.
@@ 6.2 Approach
To identify testing patterns in AI agent frameworks and agentic applications, we perform hybrid card-sorting on test functions mined from open-source repositories, as described in Section 5.3.1. Hybrid card-sorting allows us to systematically extract recurring testing patterns related to test structure, e.g., how to Arrange the test function and verification strategies and how to Assert. This process not only focuses on the existing patterns but also enables uncovering emergent patterns that have not been seen in the literature before. We provide an example of each pattern in Appendix A.
@@ 6.3 Findings
Despite the novelty of AI agent frameworks and agentic applications, 80% of their testing patterns are directly inherited from classical software engineering or ML applications, reaffirming the robustness of classical testing strategies. Table 5 summarizes the testing patterns we observed and their corresponding origins. We find that 8 out of 10 patterns have been previously reported in either ML application testing (Openja et al., 2024), or in general software engineering literature (Zhu et al., 2025; Fujita et al., 2023; Zamprogno et al., 2022; Lam et al., 2018), or in both. While ML testing research has focused more on verification patterns than on structural patterns, prior studies have still documented the use of structural patterns, e.g., test doubles (Wan et al., 2019) in ML applications. Overall, our findings suggest that practitioners are not reinventing the wheel but are strategically adapting battle-tested patterns to a new domain. The key shift is not in the patterns themselves, but in their frequency and combination.
@@
To evaluate the non-deterministic response of underlying FM in agentic applications, practitioners leverage a new state-of-the-art verification pattern, i.e., DeepEval. In general, DeepEval integrates multiple evaluation techniques, including G-Eval (Liu et al., 2023b) for assessing answer relevancy, task success, and hallucination detection, and RAGAS (Es et al., 2024) for evaluating faithfulness and precision. Both approaches leverage the LLM-as-a-judge paradigm (Zheng et al., 2023), in which a language model interprets test outputs using a set of criteria generated through Chain-of-Thought prompting (Wei et al., 2022b). Additionally, DeepEval supports user-defined logic through a rule composition mechanism based on directed acyclic graphs known as DAGMetric (Confident, 2024).
@@
We illustrate a test function that uses DeepEval in Figure 4 to explain how DeepEval helps agentic applications to tackle the uncertainty from FM. In this example, test_relevant_content_retrieved leverages DeepEval to evaluate a resume retriever agent’s output against a job description and target company. The test function invokes the G-Eval metric, which operationalizes the LLM-as-a-judge to handle non-deterministic responses. The user-provided evaluation_steps (line 18) serve as explicit evaluation criteria, instructing the judge to assess the answer’s relevancy by verifying that the correct company name is included. This approach validates the semantic correctness of the agent’s output, with the test passing if the LLM-as-a-judge’s confidence score meets the specified 0.7 threshold.
def test_relevant_content_retrieved(user_proxy: UserProxyAgent, resume_retriever_agent: ResumeRetriever, job_description: str, company: str) -> None:
message = "Here is the job description: " + job_description
chat_outcome = get_chat_outcome(user_proxy, resume_retriever_agent, message)
assert_test(
LLMTestCase(
input=message,
actual_output=chat_outcome,
context=[company]
GEval(
name="Inclusion",
evaluation_params=[
LLMTestCaseParams.INPUT,
LLMTestCaseParams.ACTUAL_OUTPUT,
threshold=0.7,
evaluation_steps=[
f"Check that the output contains the company name ’{company}’",
f"Check that the output does not contain any company name other than ’{company}’",
@@
Figure 4: Example DeepEval test case that verifies whether the retrieved output includes only the correct company name. Pattern starts by triggering assert_test on line 4, and GEval is invoked on line 11. Evaluation parameters are configured in lines 14–15, while the threshold and validation steps are defined in lines 17–20.
@@
To achieve test reproducibility, practitioners employ a hyperparameter control mechanism as a structural pattern while testing agent frameworks. We illustrate a real-world test function in Figure 5 where temperature, a hyperparameter through which randomness of FM’s output can be controlled, is set as 0 (line 6). In general, a higher temperature during generation allows FM to explore more and produce more diverse outputs (Xu et al., 2022). However, for the test function where practitioners need reproducible output every time, they set the temperature to 0 or a very low value.
def test_10_concurrent_API_calls(self):
tools = []
with open(’./data/schemas/get-headers-params.json’, r’) as f:
tools = ToolFactory.from_openapi_schema(f.read(), {})
@@
ceo = Agent(name=’CEO’, tools=tools, instructions="You are an agent that tests concurrent API calls. You must say success if the output contains headers, and error if it does not and **nothing else**.")
agency = Agency([ceo], temperature=0)
@@
result = agency.get_completion("Please call PrintHeaders tool TWICE at the same time in a single message. If any of the function outputs do not contains headers, please say error’.")
self.assertTrue(result.lower().count(’error’) == 0, agency.main_thread.thread_url)
@@
(a) Agent Framework
(b) Agentic Application
Figure 6: Co-occurrence frequency of verification patterns in the same test function in agent framework and agentic application.
@@
7.1 Motivation
A major challenge in studying agent architectures from a testing perspective is the ecosystem’s rapid and continuous evolution. Since the inception of our study in September 2024, several major updates have already reshaped the landscape: the introduction of the Model Context Protocol (MCP) for tool use in November 2024 (Hasan et al., 2025), new agent-to-agent communication primitives (A2A) in April 2025, and formal agent communication and networking protocols (ACP and ANP) in May 2025 (Ehtesham et al., 2025). To ensure our analysis is not tied to these ever evolving components and remains durable over time, we map the concrete Systems Under Test (SUTs), i.e., the specific features or modules tested in real-world agent frameworks and agentic applications, to a stable, canonical architectural component derived from established agent literature as described in Section 3.3. This canonical model enables us to systematically cover the full spectrum of architectural concerns in existing agentic systems. It also keeps our methodology extensible. For instance, as A2A emerges as a new technique (e.g., for inter-agent messaging), components using A2A can be mapped onto an existing canonical component (such as Communication Artifact).
Understanding how different components of the agent ecosystem are tested is critical for identifying whether current testing strategies adequately address component-specific risks and behaviors. For instance, Boundary Artifacts interact with external systems, which can be unreliable. This necessitates the use of testing patterns that can isolate unreliable components. On the other hand, Constitutive Entities regulate the behavior of the flow and raise exceptions when regulations are breached. Consequently, testing patterns that focus on exception handling are more relevant for constitutive entities. While RQ1 has already uncovered the types of testing patterns practitioners adopt, from novel approaches like DeepEval to the strategic adaptation of traditional less strict methods, what remains unclear is how these patterns are applied across different architectural components. Therefore, we hypothesize that testing effort is not uniform; practitioners strategically apply more intensive or specialized testing to components they perceive as riskier. By mapping the testing patterns from RQ1 to specific architectural components, we will validate this hypothesis and, more importantly, identify critical components that may be potentially under-tested.
@@ 7.3.1 Component Mapping With SUT
Belief Base Memory Persistent storage of trial history, observations, and learned knowledge (Zhang et al., 2024).
Cache Short-term semantic cache that holds recently accessed results for reuse (Bang, 2023; Zhang et al., 2025).
@@
Tokenizer Converts strings into token sequences (IDs) used internally by FMs (Rajaraman et al., 2024) and also used for knowledge and retrieving context.
@@ 7.3.2 Component-wise Test Frequency
Figure 8: Components tested in Agent Ecosystem
AI agent testing inverts the traditional ML paradigm, shifting testing effort from the non-deterministic model to the deterministic artifacts around it. Figure 8 presents the distribution of test functions across architectural components within agent frameworks and agentic applications. Our analysis shows where the testing effort is invested: Resource Artifacts receive the lion’s share of testing effort, accounting for 29.7% of tests in frameworks and 39.2% in applications. This is in stark contrast to the Plan Body (containing the FM), which receives much lower direct testing effort ($5\%$ in both cases). The disparity highlights a dramatic reallocation of engineering effort compared to traditional ML, where model testing typically commands 24–30% of the focus (Openja et al., 2024). This suggests that developers are strategically investing their limited testing resources in the components they can reliably control and verify.
Despite the Foundational Model’s strong dependency on Triggers (prompts), testing this component remains critically under-addressed, highlighting a substantial blind spot in ensuring the robustness of agents frameworks and agentic applications. As shown in Figure 8, around 1% of test functions across both frameworks and applications are dedicated to the Trigger component. This low investment is deeply concerning, as prior work has demonstrated that prompt quality, e.g., example selection, ordering, and contextual richness, significantly affects FM performance (Wu et al., 2024b). Moreover, prompt brittleness poses a growing risk as remote foundation models continue to evolve, often breaking existing prompts and degrading agent behavior unless regression testing is in place (Ma et al., 2024). The near-total absence of prompt regression testing leaves agentic systems highly vulnerable to unexpected decay in functionality, highlighting an urgent need for systematic prompt validation.
@@ 7.3.3 Component Wise Testing Patterns
For Coordination Artifacts, this means that frameworks prioritize validating the entire workflow state while applications focus on verifying final outcomes. While both employ Test Double extensively, agent frameworks show stronger adoption of Negative Testing, Snapshot Assertion, and Mock Assertion, whereas agentic applications do not use Snapshot Assertion at all and instead rely on Membership Testing approximately 2.5 times more frequently. This contrast suggests that frameworks emphasize preserving workflow integrity by capturing intermediate states, whereas applications prioritize checking artifact presence with more flexible, less rigid criteria.
@@
This contrast is further evident in testing Boundary Artifacts, where the use of Assertion-Based Testing drops from 84.6% in frameworks to 55.6% in applications, while Membership Testing increases nearly 2.9x in applications. While both employ Test Double to isolate third-party integrations, frameworks lean toward rigorous validation via Mock Assertion and Snapshot Testing, whereas agentic applications prefer lightweight, targeted checks tailored to their specific operational contexts.
The novel patterns DeepEval and Hyperparameter Control are each confined to just one or two components, explaining how their current use is highly specialized and not yet generalized across the agent ecosystem. The heatmap in Figure 9 shows that these patterns have found their initial niche use-cases. DeepEval, for instance, is used only in agentic applications for testing Observable Property and Resource Artifacts. This is its ideal niche: validating ambiguous, FM-generated artifacts, e.g., output like code, summaries, where traditional assertions fail. Likewise, Hyperparameter Control has found its primary use case in testing Roles within frameworks, directly addressing the critical need for reproducible role-wise agent behavior during development and also in Plan Body for forcing FMs to produce consistent output. While not yet in broad use, their presence in these key areas shows that novel patterns are successfully solving problems that were previously intractable, signaling their potential to expand as the agent ecosystem matures.

### v3

@@ 5.3.3 Closed Card Sorting for Components
With the SUTs identified in the previous step (Section 5.3.2), we performed a closed card-sorting to uncover the higher-level architectural components of agent frameworks and applications being tested. We considered the components described in Section 3.3 and Appendix-B as the source of components and grounded each SUT to the relevant component through the closed card sorting process.
@@
As established in the Introduction, the prevailing benchmark-driven evaluation culture often fails to capture the practical reliability of agentic systems, overlooking critical failure modes like prompt decay (Bhargava et al., 2024) and tool-use errors that are already documented in the field (Ma et al., 2024; Kokane et al., 2025). While software testing offers a path to bridge this reliability gap, the inherent non-determinism and non-reproducibility of foundation models (FMs) fundamentally challenge the application of traditional, deterministic testing techniques (Hassan et al., 2024). Hence, this RQ seeks to uncover what testing strategies actually emerge in real-world agent development, focusing on how practitioners structure tests and assert correctness despite non-deterministic outputs. The goal is to produce a foundational catalog of these real-world strategies, providing the first empirical baseline on how the open-source community is negotiating the unique quality assurance hurdles of the agentic paradigm.
@@ 6.1 Approach
To identify testing patterns in AI agent frameworks and agentic applications, we perform hybrid card-sorting on test functions mined from open-source repositories, as described in Section 5.3.1. Hybrid card-sorting allows us to systematically extract recurring testing patterns related to test structure, e.g., how to Arrange the test function and verification strategies and how to Assert. This process not only focuses on the existing patterns but also enables uncovering emergent patterns that have not been seen in the literature before. We provide an example of each pattern in Appendix A. We also compare the prevalence of each testing pattern between the top-10 popular repositories and the remaining studied repositories using Fisher’s exact test (Kim, 2017) on binary per-test-function indicators, and controlled for multiple hypothesis testing using the Benjamini–Hochberg false discovery rate (FDR) (Thissen et al., 2002) adjustment. The FDR correction limits the expected proportion of false positives among statistically significant findings, ensuring robust inference under multiple comparisons.
@@ 6.2 Findings
Despite the novelty of AI agent frameworks and agentic applications, 80% of their testing patterns are directly inherited from classical software engineering or ML applications, reaffirming the robustness of classical testing strategies. To clarify this landscape, we distinguish between traditional adaptations, which extend classical paradigms to structure and verify the agents, and non-determinism-specific innovations, which emerge uniquely to address the non-deterministic characteristics of LLM-driven systems in Table 6. We find that 8 out of 10 patterns have been previously reported in either ML application testing (Openja et al., 2024), or in general software engineering literature (Zhu et al., 2025; Fujita et al., 2023; Zamprogno et al., 2022; Lam et al., 2018), or in both. While ML testing research has focused more on verification patterns than on structural patterns, prior studies have still documented the use of structural patterns, e.g., test doubles (Wan et al., 2019) in ML applications. As the majority of observed patterns fall into the former category, this indicates strong continuity with established testing traditions. Overall, our findings suggest that practitioners are not reinventing the wheel but are strategically adapting battle-tested patterns to a new domain. The key shift is not in the patterns themselves, but in their frequency and combination.
@@
To evaluate the non-deterministic response of underlying FM in agentic applications, practitioners leverage a new state-of-the-art verification pattern, i.e., DeepEval. This pattern represents a necessary non-determinism-specific innovation, as classical assertion mechanisms (e.g., exact matching) cannot accommodate the semantic variability inherent in LLM outputs. Unlike conventional test oracles, DeepEval is designed to orchestrate the evaluation of interactive and non-deterministic model behavior (e.g., creativity) under task-specific criteria, making it particularly suitable for agentic systems (Kim and Oh, 2025; Mohammadi et al., 2025). In general, DeepEval integrates multiple evaluation techniques, including G-Eval (Liu et al., 2023b) for assessing answer relevancy, task success, and hallucination detection, and RAGAS (Es et al., 2024) for evaluating faithfulness and precision. Both approaches leverage the LLM-as-a-judge paradigm (Zheng et al., 2023), in which a language model interprets test outputs using a set of criteria generated through Chain-of-Thought prompting (Wei et al., 2022b). Additionally, DeepEval supports user-defined logic through a rule composition mechanism based on directed acyclic graphs known as DAGMetric (Confident, 2024).
@@
We illustrate a test function that uses DeepEval in Figure 5 to explain how DeepEval helps agentic applications to tackle the uncertainty from FM. In this example, test_relevant_content_retrieved leverages DeepEval to evaluate a resume retriever agent’s output against a job description and target company. By assessing whether the agent output satisfies a semantic objective rather than matching a fixed string, this pattern directly targets the kind of non-deterministic behavior that arises in agent execution. The test function invokes the G-Eval metric, which operationalizes the LLM-as-a-judge to handle non-deterministic responses. The user-provided evaluation_steps (line 18) serve as explicit evaluation criteria, instructing the judge to assess the answer’s relevancy by verifying that the correct company name is included. This approach validates the semantic correctness of the agent’s output, with the test passing if the LLM-as-a-judge’s confidence score meets the specified 0.7 threshold.
1def test_relevant_content_retrieved(
2 user_proxy: UserProxyAgent,
3 resume_retriever_agent: ResumeRetriever,
4 job_description: str,
5 company: str
6) -> None:
7 message = "Here is the job description: " + job_description
9 assert_test(
10 LLMTestCase(
11 input=message,
12 actual_output=chat_outcome,
13 context=[company]
14 ),
15 [
16 GEval(
17 name="Inclusion",
18 evaluation_params=[
19 LLMTestCaseParams.INPUT,
20 LLMTestCaseParams.ACTUAL_OUTPUT,
21 ],
22 threshold=0.7,
23 evaluation_steps=[
24 f"Check that the output contains the company name ’{company}’",
26 ],
27 )
28 ],
29 )
@@ 16 GEval(
Figure 5: Example DeepEval test case that verifies whether the retrieved output includes only the correct company name. Pattern starts by triggering assert_test, and GEval is invoked subsequently. Evaluation parameters, threshold, and validation steps are configured in the highlighted lines.
@@
To achieve test reproducibility, practitioners employ a hyperparameter control mechanism as a structural pattern while testing agent frameworks. We illustrate a real-world test function in Figure 6 where “temperature”, a hyperparameter through which randomness of FM’s output can be controlled, is set as 0 (line 6). In general, a higher temperature during generation allows FM to explore more and produce more diverse outputs (Xu et al., 2022). However, for the test function where practitioners need reproducible output every time, they set the temperature to 0 or a very low value. Unlike traditional configuration management, this pattern specifically targets the stochastic nature of the underlying model, marking it as an innovation designed to enforce determinism in an inherently probabilistic environment.
1def test_10_concurrent_API_calls(self):
2 tools = []
3 with open(’./data/schemas/get-headers-params.json’, ’r’) as f:
4 tools = ToolFactory.from_openapi_schema(f.read(), {})
5 ceo = Agent(
6 name=’CEO’,
7 tools=tools,
@@
8 instructions="You are an agent that tests concurrent API calls. You must say ’success’ if the output contains headers, and ’error’ if it does not and **nothing else**."
9 )
10 agency = Agency([ceo], temperature=0)
11 result = agency.get_completion(
@@
12 "Please call PrintHeaders tool TWICE at the same time in a single message. "
13 "If any of the function outputs does not contain headers, please say ’error’."
14 )
15 self.assertTrue(
16 result.lower().count(’error’) == 0,
17 agency.main_thread.thread_url
18 )
@@
(a) Agent Frameworks
(b) Agentic Applications
Figure 7: Co-occurrence frequency of verification patterns in the same test function in agent frameworks and agentic applications.
Highly popular agent frameworks do not adopt testing patterns differently from the rest of the agent ecosystem. To examine this, we compare the ten most popular agent frameworks (ranked by GitHub star count) against all remaining repositories using Fisher’s Exact Test for each testing pattern, and present our results in Table 7. The analysis shows that, after correcting for multiple comparisons using FDR, only Mock Assertion-based testing is significantly more prevalent in top-10 frameworks (22.9% vs. 9.1%, $p=0.0004$, $p_{\text{adj}}=0.0049$). In contrast, all other testing patterns exhibit no statistically significant differences. Similarly, for agentic applications, we do not find any pattern significantly dominating in the top-10 agentic applications.
Table 7: Comparison of testing pattern prevalence between top-10 and other repositories. Statistical significance is evaluated using p-value and adjusted p-value (FDR correction).
Testing Pattern Overall (%) Top10 (%) Others (%) p-value p_adj
Mock Assertion 14.9 22.9 9.1 0.0004 0.0049
Test Double 24.4 29.9 20.5 0.0389 0.1428
Parameterized Testing 28.7 33.8 25.0 0.0658 0.1449
Assertion Based Testing 70.0 72.6 68.2 0.3644 0.6293
Functional 2.9 5.1 1.4 0.0577 0.1449
Value Range Analysis 3.4 4.5 2.7 0.4004 0.6293
DeepEval 0.0 0.0 0.0 1.0000 1.0000
Hyperparameter Control 0.5 0.0 0.9 0.5126 0.7049
Membership Testing 15.9 14.7 16.8 0.6686 0.7355
Negative Test 24.4 22.9 25.5 0.6272 0.7355
Snapshot Testing 5.3 1.9 7.7 0.0176 0.0972
@@
Identifying testing patterns alone is insufficient without understanding where they are applied within agent architectures. FM-based agents comprise heterogeneous components with fundamentally different failure modes, ranging from deterministic infrastructure to non-deterministic decision logic. Consequently, uniform testing effort across components would be both impractical and suboptimal.
This RQ examines how practitioners distribute testing effort across architectural components and whether testing strategies align with component-specific risks. By mapping observed tests to a stable canonical architecture, we aim to reveal systematic testing blind spots and prioritization biases that would otherwise remain obscured by framework-specific implementations.
@@ 7.2.1 Component Mapping With SUT
Belief Base Memory Persistent storage of trial history, observations, and learned knowledge (Zhang et al., 2025a).
Cache Short-term semantic cache that holds recently accessed results for reuse (Bang, 2023; Zhang et al., 2025b).
@@
Tokenizer Converts strings into token sequences (IDs) used internally by FMs (Schmidt et al., 2024) and also used for knowledge and retrieving context.
@@ 7.2.2 Component-wise Test Frequency
(a) Distribution of test function count across architectural components.
(b) Distribution of test Lines of Code (LoC) across architectural components.
Figure 9: Components tested in the agent ecosystem.
AI agent testing inverts the traditional ML paradigm, shifting testing effort from the non-deterministic model to the deterministic artifacts around it. Figure 9 presents the distribution of test functions across architectural components within agent frameworks and agentic applications. To quantify the testing effort, we triangulate our findings using two complementary metrics: (i) the raw count of test functions, a standard proxy in ML testing research (Openja et al., 2024), and (ii) the Lines of Code (LoC) within those tests, a widely accepted measure of implementation effort in software engineering (Labuschagne et al., 2017). Together, these metrics capture both how often components are tested and how much coding effort is invested in testing them. Our analysis shows where the testing effort is invested: Resource Artifacts receive the lion’s share of testing effort, accounting for 29.7% (tested in 112 functions out of 377 sampled functions) of test functions in frameworks (27.8% LoC, consuming 1,209 lines out of 4,354 lines of test code) and 39.2% (tested in 149 functions out of 382 sampled functions) in applications (35.2% LoC, 1480 out of 4,197 LoC). This is in stark contrast to the Plan Body (containing the FM), which receives much lower direct testing effort accounting for 4.8% (tested in 18 functions out of 377 sampled functions) of test functions in frameworks (4.7% LoC, 203 out of 4,354 LoC) and 5.0% (tested in 19 functions out of 382 sampled functions) in applications (5.2% LoC, 220 out of 4,197 LoC). The disparity highlights a dramatic reallocation of engineering effort compared to traditional ML, where model testing typically consumes 24–30% of the effort (Openja et al., 2024). This suggests that developers are strategically allocating their limited testing resources to components they can reliably control and verify.
Despite the Foundation Model’s strong dependency on Triggers (prompts), testing this component remains critically under-addressed, highlighting a substantial blind spot in ensuring the robustness of agents frameworks and agentic applications. As shown in Figure 9, around 1% of test functions across both frameworks and applications are dedicated to the Trigger component. This low investment is deeply concerning, as prior work has demonstrated that prompt quality, e.g., example selection, ordering, and contextual richness, significantly affects FM performance (Wu et al., 2024b). Moreover, prompt brittleness poses a growing risk as remote foundation models continue to evolve, often breaking existing prompts and degrading agent behavior unless regression testing is in place (Ma et al., 2024). The near-total absence of prompt regression testing leaves agentic systems highly vulnerable to unexpected decay in functionality, highlighting an urgent need for systematic prompt validation.
@@ 7.2.3 Component Wise Testing Patterns
For Coordination Artifacts, this means that frameworks prioritize validating the entire workflow state while applications focus on verifying final outcomes. Both agent frameworks and agentic applications rely heavily on Test Double for testing Coordination Artifact. Agent frameworks complement this pattern by using Negative Testing, Snapshot Assertion, and Mock Assertion to a markedly greater extent, reflecting a focus on validating intermediate workflow states. By contrast, agentic applications entirely forgo Snapshot Assertion and instead rely on Membership Testing, adopting it approximately 2.5 times more frequently to accommodate flexible, outcome-oriented validation for this component. This contrast suggests that frameworks emphasize preserving workflow integrity by capturing intermediate states, whereas applications prioritize checking artifact presence with more flexible, less rigid criteria.
@@
This contrast is further evident in testing Boundary Artifacts, where the use of Assertion-Based Testing drops from 84.6% in frameworks to 55.6% in applications, while Membership Testing increases nearly 2.9x in applications. Both agent frameworks and agentic applications rely on Test Double to isolate third-party integrations. Agent frameworks complement this pattern with greater use of Mock Assertion and Snapshot Testing, reflecting an emphasis on stricter validation of integration behavior. In contrast, agentic applications de-emphasize these techniques and instead favor lighter, context-specific checks aligned with their operational needs.
The novel patterns DeepEval and Hyperparameter Control are each confined to just one or two components, explaining how their current use is highly specialized and not yet generalized across the agent ecosystem. The heatmap in Figure 10 shows that these patterns have found their initial niche use-cases. DeepEval, for instance, is used only in agentic applications for testing Observable Property and Resource Artifacts. This is its ideal niche: validating ambiguous, FM-generated artifacts, e.g., output like code, summaries, where traditional assertions fail. Likewise, Hyperparameter Control is primarily used to test Roles within frameworks, where practitioners use hyperparameters (e.g., temperature) to ensure reproducible role-wise agent behavior during testing. It is also applied in the Plan Body to obtain more consistent FM outputs. While not yet in broad use ( less than 1% in both agent frameworks and agentic application), their presence in these key areas shows that novel patterns are successfully solving problems that were previously intractable, signaling their potential to expand as the agent ecosystem matures.